OpenStudy (anonymous):

IF YOU HELP YOU GET A MEDAL!!! You are making a square quilt with the design shown below. Find the side length x of the inner square that would make its area equal to 30% of the total area of the quilt. Round to the nearest tenth of a foot. 3 ft 3.3 ft 4.2 ft 12 ft

OpenStudy (anonymous):

here is the photo

OpenStudy (saifoo.khan):

Find the total area, then find it's 30%

OpenStudy (anonymous):

3.3

OpenStudy (anonymous):

30% of 36 = 10.8 \[\sqrt{10.8} \approx 3.3\]

OpenStudy (saifoo.khan):

Area of the bigger square = 6*6 = 36 units^2 Now find it's 30% \[\frac{30}{100} \times 36\] \[10.8 units^2\] Now area = l^2 10.8 = l^2 l = sqrt(10.8) l = 3.28 units. So x = 3.28